OpenStudy (anonymous):

Find all the critical numberes of x2^(-x^2)

OpenStudy (anonymous):

Do you mean X to the power or 2 to the power? @calc_nofori

OpenStudy (anonymous):

\[x2^{-x ^{^{2}}}\]

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} f(x)=x\times 2^(-x^2)\hspace{.33em}\\~\\ \end{align}}\) find \(\large \color{black}{\begin{align} f(x)'\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

i just dont know how. i know the multiplication rule but what confuses me is the derivative of 2^-x^2

OpenStudy (mathmath333):

this is ur number right \(\large \color{black}{\begin{align} f(x)=x\times 2^{-x^2}\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{d}{dx}a^x=a^x \times \log_{e}a\] Apply chain rule

OpenStudy (anonymous):

\[2^{-x ^{2}}\log_{e} 2\]

OpenStudy (anonymous):

then itd be times -2x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you may also combine the 2 in -2x and the 2 in 2^(-x^2) by adding the exponents

OpenStudy (anonymous):

and then i'd set that to 0 to get the critical numbers?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} f(x)=x\times 2^{-x^2}\hspace{.33em}\\~\\ f'(x)=x\times \left(2^{-x^2}\right)'+x'\times \left(2^{-x^2}\right)\hspace{.33em}\\~\\ f'(x)=x\times \dfrac{d}{d(-x^2)}\left(2^{-x^2}\right) \dfrac{d-x^2}{dx}+\left(2^{-x^2}\right)\hspace{.33em}\\~\\ f'(x)=x\times \log 2\times \left(2^{-x^2}\right) (-2x)+\left(2^{-x^2}\right)\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

yep f'(x)=0 and find the values of x, they will be the points on the curve for which you have local maxima or minima

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} f'(x)=\left(2^{-x^2}\right)(1-x^2 \log 4)\hspace{.33em}\\~\\ \left(2^{-x^2}\right)(1-x^2 \log 4)=0\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

it'd be 0 and then for the second part, i wouldnt know how to do it properly. I know the answer is suppose to be \[1/\sqrt{2\ln2}\]

OpenStudy (anonymous):

but i dont get that answer

OpenStudy (mathmath333):

\(\large \co\(\large \color{black}{\begin{align} f'(x)&=\left(2^{-x^2}\right)(1-x^2 \log 4)\hspace{.33em}\\~\\ \left(2^{-x^2}\right)(1-x^2 \log 4)&=0\hspace{.33em}\\~\\ \left(2^{-x^2}\right)&=0\hspace{.33em}\\~\\ \text{or}\hspace{.33em}\\~\\ (1-x^2 \log 4)&=0\hspace{.33em}\\~\\ \left(2^{-x^2}\right)&=2^0\hspace{.33em}\\~\\ \text{or}\hspace{.33em}\\~\\ x^2 \log 4&=1\hspace{.33em}\\~\\ -x^2&=0\hspace{.33em}\\~\\ \text{or}\hspace{.33em}\\~\\ x^2& =\dfrac{1}{\log 4}\hspace{.33em}\\~\\ x&=0\hspace{.33em}\\~\\ \text{or}\hspace{.33em}\\~\\ x&=\pm\sqrt{\dfrac{1}{2\log 2}}\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

forget /ignore the red

OpenStudy (anonymous):

@mathmath333 wouldn't 2^0 be equal to 1 and not 0

OpenStudy (mathmath333):

oh yep let me correct

OpenStudy (anonymous):

how do you get it to be \[\sqrt{\frac{ 1 }{ 2\log2 }}\]

OpenStudy (anonymous):

what do you mean @calc_nofori The solution of the equation is x=1/sqrt(2log2). That's about it, if you want to find the point, you'll have to put the value of x in your f(x), that will give you your 'y'.

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \left(2^{-x^2}\right)&=0\hspace{.33em}\\~\\ \log \left(2^{-x^2}\right)&=\log 0\hspace{.33em}\\~\\ -x^2\log \left(2\right)&=-\infty \hspace{.33em}\\~\\ -x^2&=-\infty \hspace{.33em}\\~\\ x&=\infty \end{align}}\)